OpenStudy (anonymous):

Please and thank you! :) Fun Land theme park has a total of 140 males and females who are employed at the park. If the number of males is 20 more than twice the number of females, how many males are employed at the park?

OpenStudy (anonymous):

what is 140 times 2?

OpenStudy (anonymous):

280 right?

OpenStudy (anonymous):

yes! what is 280+20?

OpenStudy (anonymous):

300

OpenStudy (anonymous):

and that right there is your answer.

OpenStudy (anonymous):

40 60 100 110 mutiple choices..........

OpenStudy (anonymous):

oh pellet I read the question wrong.

OpenStudy (callisto):

Let x be the number of female. no of male = 2x +20 x +(2x+20) =140 Can you solve it from here?

OpenStudy (anonymous):

A lil confusing :\ im not sure on how to sovle it

OpenStudy (callisto):

First let x be the number of female Since 'number of males is 20 more than twice the number of females' number of male = 20 + 2x Understand this part?

OpenStudy (callisto):

Sorry I have to go.. so i leave you some explanations here for the answer.. Then, from 'has a total of 140 males and females who are employed at the park' you know no. of male + no of female = 140 So (2x+20) + x =140 Then solve for x After that, find the no of male by putting the value of x into 2x+20

OpenStudy (anonymous):

Is it 100?

OpenStudy (callisto):

Steps for solving x : (2x+20) + x =140 2x + x +20 = 140 3x =120 x = 120 /3 And yes, your answer should be correct